# Dual Ascent Diffusion for Inverse Problems

### v1

@@ 1 Introduction
We are interested in solving inverse problems, where an unknown image or signal $\mathbf{x}$ is estimated from noisy and corrupted observations $\mathbf{y}$. These types of problems arise across science and engineering, for example, in image restoration [24], astrophysics [1], medical imaging [35], protein structure determination [14, 25, 22], among other domains. In all cases, a linear or nonlinear function $\mathcal{A}\left(\cdot\right)$ models a domain-specific image formation process. Although the likelihood of observations $p\left(\mathbf{y}|\mathbf{x}\right)$ depends on the statistical model of the noise in the observations, closed-form expressions exist for specific cases. For example, the image formation model for zero-mean Gaussian i.i.d. noise with variance $\sigma^{2}$ is $\mathbf{y}=\mathcal{A}\left(\mathbf{x}\right)+\mathcal{N}\left(\mathbf{0},\sigma^{2}\right)$ and its log-likelihood $\log p\left(\mathbf{y}|\mathbf{x}\right)=\frac{-1}{2\sigma^{2}}\left\|\mathbf{y}-\mathcal{A}\left(\mathbf{x}\right)\right\|_{2}^{2}$, up to an additive term that does not depend on $\mathbf{x}$. Given the likelihood of the observations and a prior $p\left(\mathbf{x}\right)$, inverse problem solvers aim at either maximizing or sampling from the posterior $p(\mathbf{x}|\mathbf{y})\propto p(\mathbf{y}|\mathbf{x})p(\mathbf{x})$ in a Bayesian framework. Most inverse problems are ill-posed, making the prior a crucial component of the solution-finding process.
@@
Maximum-a-posteriori (MAP) approaches aim at maximizing the posterior to find the most likely solution given a prior. Early approaches used “hand-crafted” priors to promote smoothness, piece-wise constancy via Total Variation [29, 3], or sparsity in a transform domain [12], while most modern approaches use some form of neural network [40]. The plug-n-play (PnP) approach [37], for example implemented by the Alternating Direction Method of Moments (ADMM) [4] algorithm, is a popular and versatile framework to solve MAP problems by leveraging (Gaussian) denoisers as priors. MAP finds the single, most likely solution to an inverse problem. However, oftentimes one is interested in sampling from the posterior of all feasible solutions. For this reason, many recent works [18, 33, 6, 7, 32, 20, 21, 38, 15, 26, 39] focus on posterior sampling using powerful pretrained diffusion models as priors, as surveyed in [9]. While these recent diffusion posterior sampling methods show great promise, they are all fundamentally limited by the optimization framework that is used to combine the likelihood of the image formation model and the prior during optimization.
@@
In this work, we do not aim to develop a method that provably samples from the posterior, but instead focus on deriving an optimization strategy that accurately and efficiently solves the MAP problem using a prior given by a pretrained diffusion model and a dual ascent–based optimization framework inspired by ADMM [4]. Our approach, dubbed DDiff, is faster and shown to achieve better reconstruction quality compared with the state of the art for image restoration problems, including single-image super resolution, inpainting, deblurring, phase retrieval, and high-dynamic range imaging. Moreover, DDiff is more robust to high levels of measurement noise, and our reconstructions more faithfully model the observations by exhibiting a lower likelihood than existing methods. The latter is important because the log-likelihood of the observations $\textrm{log}\,p\left(\mathbf{y}|\mathbf{x}\right)$ is an indicator for the level of hallucination a generative prior, such as a diffusion model, introduces when computing a solution.
@@ 2.1 Maximum-a-Posteriori Solutions
$\mathbf{x}_{\textrm{MAP}}=\operatorname*{argmin}_{\mathbf{x}}-\left(\log p\left(\mathbf{y}|\mathbf{x}\right)+\log p\left(\mathbf{x}\right)\right)=\frac{1}{2\sigma^{2}}\left\|\mathbf{y}-\mathcal{A}\left(\mathbf{x}\right)\right\|_{2}^{2}-\log p\left(\mathbf{x}\right).$ (1)
The alternating direction method of multipliers (ADMM) [4] is a common approach to solving the MAP problem. ADMM attempts to blend the benefits of dual decomposition and augmented Lagrangian methods for constrained optimization. For this purpose, a slack variable $\mathbf{z}$ is introduced to split the objective function in Eq. 1 into a data fidelity term $-\log p\left(\mathbf{y}|\mathbf{x}\right)$ and the log-prior term $-\log p\left(\mathbf{z}\right)$, subject to $\mathbf{x}=\mathbf{z}$. ADMM then forms the Augmented Lagrangian of the split formulation as
$L_{\rho}\left(\mathbf{x},\mathbf{z},\mathbf{u}\right)=\frac{1}{2\sigma^{2}}\left\|\mathbf{y}-\mathcal{A}\left(\mathbf{x}\right)\right\|_{2}^{2}-\log p\left(\mathbf{z}\right)+\frac{\rho}{2}\left\|\mathbf{x}-\mathbf{z}+\mathbf{u}\right\|_{2}^{2}-\frac{\rho}{2}\left\|\mathbf{u}\right\|_{2}^{2},$ (2)
@@
$\mathbf{z}$ $\leftarrow\operatorname*{argmin}_{\mathbf{z}}-\log p\left(\mathbf{z}\right)+\frac{\rho}{2}\|\mathbf{x}-\mathbf{z}+\mathbf{u}\|_{2}^{2}=\mathcal{D}\left(\mathbf{x}+\mathbf{u},\tilde{\sigma}^{2}=\frac{1}{\rho}\right)$ (4)
@@
2.2 Diffusion Models and Diffusion Posterior Sampling
@@
$\mathbf{x}_{t-1}=\frac{1}{\sqrt{{\alpha}_{t}}}\left(\mathbf{x}_{t}+\left(1-{\alpha}_{t}\right)\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\right)+\sqrt{1-{\alpha}_{t}}\,\epsilon,\quad\epsilon\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}\right)$ (6)
for $t=T\ldots 1$ to generate an image $\mathbf{x}_{0}$. The reverse diffusion process [2, 16] approximates the inverse trajectories of a corresponding forward diffusion $\mathbf{x}_{t}=\sqrt{\bar{\alpha}}_{t}\,\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon$, $\epsilon\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}\right)$. Here, we adopt the variance-preserving form of forward and reverse diffusion [17]. The factors $\alpha_{t}$ and $\bar{\alpha}_{t}=\prod_{s=0}^{t}\alpha_{s}$ are derived from the noise schedule of the diffusion model [17]. Importantly, the score network $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$, defined by parameters $\theta$, is a neural network that approximates the score function $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ [34]. This network is learned from training data in the diffusion model pretraining stage. To sample more efficiently, Denoising Diffusion Implicit Models (DDIM) [31] converts Eq. 6 into a non-Markovian process:
$\mathbf{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{\mathbf{x}_{t}+(1-\bar{\alpha}_{t})\mathbf{s}_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)-\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\,\left(\sqrt{1-\bar{\alpha}_{t}}\cdot\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\right)\,+\sigma_{t}\epsilon,$ (7)
@@
In a posterior sampling problem, we aim at sampling from the posterior $p(\mathbf{x}|\mathbf{y})$. For this purpose, many recent methods follow the approach described above, using the posterior score $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}|\mathbf{y})=\nabla_{\mathbf{x}}\log p(\mathbf{y}|\mathbf{x}_{t}=\mathbf{x})+\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ instead of $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$. The second term is equivalent to the unconditional score of the pretrained diffusion model, but the challenge lies in the first, i.e., the conditional score term. The conditional probability $p(\mathbf{y}|\mathbf{x}_{t})$ can be written as a conditional expectation $\mathbb{E}_{\mathbf{x}_{0}\sim p(\mathbf{x}_{0}|\mathbf{x}_{t})}[p(\mathbf{y}|\mathbf{x}_{0})]$, but approximating this expectation with Monte Carlo samples is computationally intractable (see Chung et al. [7], for example). For this reason, existing diffusion posterior sampling methods approximate this conditional distribution with a Dirac delta distribution concentrated on $\mathbf{x}_{t}$ [18] or on $\mathbb{E}_{\mathbf{x}_{0}\sim p(\mathbf{x}_{0}|\mathbf{x}_{t})}[\mathbf{x}_{0}]$ [7]. A number of other diffusion posterior sample methods have been proposed [9], each providing a different approximation for the conditional expectation. Most recently, Zhang et al. [39] introduced DAPS (Decoupled Annealing Posterior Sampling), a two-step iterative approach that mitigates the accumulation of errors along the sampling trajectory through an Markov chain Monte Carlo-based equilibration step at $t=0$. DAPS demonstrated state-of-the-art results on both linear and nonlinear inverse problems, but remains limited by the necessity to approximate the conditional probability $p(\mathbf{x}_{0}|\mathbf{x}_{t})$ and by the finiteness of the number of MCMC steps.
@@ 3.2 Dual Ascent Diffusion (DDiff)
At its core, the ADMM method iterates over 3 steps. The $\mathbf{x}$-update corresponds to a data matching step (Eq. 3), the $\mathbf{z}$-update can be seen as a denoising step (Eq. 4) and the dual update stems from the introduction of the dual variable $\mathbf{u}$ for the constraint $\mathbf{x}=\mathbf{z}$ (Eq. 5). While the $\mathbf{x}$ and dual updates are straightforwardly derived from the original ADMM framework, our methodological contribution primarily consists in showing that a pretrained diffusion model can be used more efficiently in the $\mathbf{z}$-update than the naive approach (Eq. 8).
@@
Combined in an iterative fashion, these three steps define our method, described with pseudo-code in Algorithm 1.
Figure 1: Overview of DDiff. This example illustrates the motion deblurring task. Our method alternates between three steps ($\mathbf{x}$-update, $\mathbf{z}$-update, dual update). On the right, we show the evolution of $\mathbf{x}$, $\mathbf{z}$, and $\mathbf{u}$ throughout optimization.
@@ 3.3 Comparison to Other Diffusion-Based Variable Splitting Methods
Prior methods, such as DiffPIR [42] and DCDP [23], are MAP-based optimization methods, specifically built upon the half-quadratic splitting (HQS) method, which lacks dual variables in its formulation. Our work naturally extends existing frameworks, transforming them from HQS-style approaches to those using dual variables, including ADMM. By incorporating Lagrange multipliers that accumulate constraint violations across the iterations, DDiff offers improved empirical performance on challenging inverse problems where measurement consistency is crucial. In particular, we note that removing the dual update from DDiff would exactly emulate DiffPIR [42] (with the right choice of $\sigma_{t}$ and $\gamma_{t}$, see supplements).
@@
2:  for $t=T-1$ to 00 do
3:     $\mathbf{z}\leftarrow\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\mathbf{x}_{t}+(1-\bar{\alpha}_{t})\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\right)$ Denoising step (Eq. 10)
4:     $\mathbf{x}\leftarrow\mathbf{z}-\mathbf{u}-\gamma_{t}\nabla_{\mathbf{v}=\mathbf{z}-\mathbf{u}}\|\mathbf{y}-\mathcal{A}(\mathbf{v})\|^{2}$ Measurement step (Eq. 9)
7:     $\mathbf{x}_{t-1}\leftarrow\sqrt{\bar{\alpha}_{t-1}}\cdot\mathbf{x}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\hat{\epsilon}+\sigma_{t}\epsilon+\sqrt{\bar{\alpha}_{t-1}}\cdot\mathbf{u}$ Reverse diffusion (Eq. 11)
8:     $\mathbf{u}\leftarrow\mathbf{u}+\mathbf{x}-\mathbf{z}$Dual update (Eq. 5)
9:  end for
10:  return  $\mathbf{x}_{0}$

### v2

@@ 1 Introduction
We are interested in solving inverse problems, where an unknown image or signal $\mathbf{x}$ is estimated from noisy and corrupted observations $\mathbf{y}$. These types of problems arise across science and engineering, for example, in image restoration [25], astrophysics [1], medical imaging [36], protein structure determination [14, 26, 23], among other domains. In all cases, a linear or nonlinear function $\mathcal{A}\left(\cdot\right)$ models a domain-specific image formation process. Although the likelihood of observations $p\left(\mathbf{y}|\mathbf{x}\right)$ depends on the statistical model of the noise in the observations, closed-form expressions exist for specific cases. For example, the image formation model for zero-mean Gaussian i.i.d. noise with variance $\sigma^{2}$ is $\mathbf{y}=\mathcal{A}\left(\mathbf{x}\right)+\mathcal{N}\left(\mathbf{0},\sigma^{2}\right)$ and its log-likelihood is $\log p\left(\mathbf{y}|\mathbf{x}\right)=\frac{-1}{2\sigma^{2}}\left\|\mathbf{y}-\mathcal{A}\left(\mathbf{x}\right)\right\|_{2}^{2}$, up to an additive term that does not depend on $\mathbf{x}$. Given the likelihood of the observations and a prior $p\left(\mathbf{x}\right)$, inverse problem solvers aim at either maximizing or sampling from the posterior $p(\mathbf{x}|\mathbf{y})\propto p(\mathbf{y}|\mathbf{x})p(\mathbf{x})$ in a Bayesian framework. Most inverse problems are ill-posed, making the prior a crucial component of the solution-finding process.
@@
Maximum-a-posteriori (MAP) approaches aim at maximizing the posterior to find the most likely solution given a prior. Early approaches used “hand-crafted” priors to promote smoothness, piece-wise constancy via Total Variation [31, 3], or sparsity in a transform domain [12], while most modern approaches use some form of neural network [44]. The plug-n-play (PnP) approach [39], for example implemented by the Alternating Direction Method of Multipliers (ADMM) [4] algorithm, is a popular and versatile framework to solve MAP problems by leveraging (Gaussian) denoisers as priors. MAP finds the single, most likely solution to an inverse problem. However, oftentimes one is interested in sampling from the posterior of all feasible solutions. For this reason, many recent works [19, 37, 6, 7, 34, 22, 21, 41, 15, 27, 43] focus on posterior sampling using powerful pretrained diffusion models as priors, as surveyed in [9]. While these recent diffusion posterior sampling methods show great promise, they are all fundamentally limited by the optimization framework that is used to combine the likelihood of the image formation model and the prior during optimization.
@@
In this work, we do not aim to develop a method that provably samples from the posterior, but instead focus on deriving an optimization strategy that accurately and efficiently solves the MAP problem using a prior given by a pretrained diffusion model and a dual ascent–based optimization framework inspired by ADMM [4]. Our approach, dubbed DDiff, is faster and shown to achieve better reconstruction quality compared with the state of the art for image restoration problems, including single-image super resolution, inpainting, deblurring, phase retrieval, and high-dynamic range imaging. Moreover, DDiff is more robust to high levels of measurement noise, and our reconstructions more faithfully model the observations by exhibiting closer-to-zero residuals than existing methods. The latter is important because the log-likelihood of the observations $\textrm{log}\,p\left(\mathbf{y}|\mathbf{x}\right)$ is an indicator for the level of hallucination a generative prior, such as a diffusion model, introduces when computing a solution. Notably, we establish the first fixed-point convergence proof for a diffusion-based posterior optimization method and extend our framework to a latent setting for efficient inference in compact feature spaces.
@@ 2.1 Maximum-a-Posteriori Solutions
$\mathbf{x}_{\mathrm{MAP}}$ $=\operatorname*{argmin}_{\mathbf{x}}\;-\big(\log p(\mathbf{y}\mid\mathbf{x})+\log p(\mathbf{x})\big)$ (1)
$=\frac{1}{2\sigma^{2}}\left\|\mathbf{y}-\mathcal{A}(\mathbf{x})\right\|_{2}^{2}\;-\;\log p(\mathbf{x})$
The alternating direction method of multipliers (ADMM) [4] is a common approach to solving the MAP problem. ADMM attempts to blend the benefits of dual decomposition and augmented Lagrangian methods for constrained optimization. For this purpose, a slack variable $\mathbf{z}$ is introduced to split the objective function in Eq. 1 into a data fidelity term $\log p\left(\mathbf{y}|\mathbf{x}\right)$ and the log-prior term $\log p\left(\mathbf{z}\right)$, subject to $\mathbf{x}=\mathbf{z}$. ADMM then forms the Augmented Lagrangian of the split formulation as
$L_{\rho}\left(\mathbf{x},\mathbf{z},\mathbf{u}\right)$ $=\frac{1}{2\sigma^{2}}\left\|\mathbf{y}-\mathcal{A}\left(\mathbf{x}\right)\right\|_{2}^{2}-\log p\left(\mathbf{z}\right)$ (2)
$\quad+\frac{\rho}{2}\left\|\mathbf{x}-\mathbf{z}+\mathbf{u}\right\|_{2}^{2}-\frac{\rho}{2}\left\|\mathbf{u}\right\|_{2}^{2},$
@@
$\mathbf{z}$ $\leftarrow\operatorname*{argmin}_{\mathbf{z}}-\log p\left(\mathbf{z}\right)+\frac{\rho}{2}\|\mathbf{x}-\mathbf{z}+\mathbf{u}\|_{2}^{2}$ (4)
$=\mathcal{D}\left(\mathbf{x}+\mathbf{u},\tilde{\sigma}^{2}=\frac{1}{\rho}\right)$ (5)
@@
Figure 1: Overview of DDiff. This example illustrates motion deblurring. DDiff alternates between three updates ($\mathbf{x}$-, $\mathbf{z}$-, and dual updates) within each ADMM iteration. As iterations proceed, the primal variables $\mathbf{x}$ and $\mathbf{z}$ progressively align, while the dual variable $\mathbf{u}$ diminishes toward zero, indicating convergence to a fixed point $(\mathbf{x}^{\star},\mathbf{z}^{\star},\mathbf{u}^{\star})$. This evolution visually demonstrates the fixed-point convergence behavior analyzed in Appendix A. Here, the ADMM iteration index and diffusion timestep are set equal by design.
2.2 Diffusion Models and Posterior Sampling
@@
$\mathbf{x}_{t-1}=\frac{1}{\sqrt{{\alpha}_{t}}}\left(\mathbf{x}_{t}+\left(1-{\alpha}_{t}\right)\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\right)+\sqrt{1-{\alpha}_{t}}\,\epsilon$ (7)
for $t=T,\ldots,1$ to generate an image $\mathbf{x}_{0}$, where $\epsilon\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}\right)$. The reverse diffusion process [2, 17] approximates the inverse trajectories of a corresponding forward diffusion $\mathbf{x}_{t}=\sqrt{\bar{\alpha}}_{t}\,\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon$, $\epsilon\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}\right)$. Here, we adopt the variance-preserving form of forward and reverse diffusion [18]. The factors $\alpha_{t}$ and $\bar{\alpha}_{t}=\prod_{s=0}^{t}\alpha_{s}$ are derived from the noise schedule of the diffusion model [18]. Importantly, the score network $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$, defined by parameters $\theta$, is a neural network that approximates the score function $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ [35]. This network is learned from training data in the diffusion model pretraining stage. To sample more efficiently, Denoising Diffusion Implicit Models (DDIM) [33] provides an alternative non-Markovian reverse parameterization of the diffusion process, replacing the Markovian formulation in Eq. 7:
$\mathbf{x}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{\mathbf{x}_{t}+(1-\bar{\alpha}_{t})\mathbf{s}_{\theta}(\mathbf{x}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)$ (8)
$\quad-\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\,\left(\sqrt{1-\bar{\alpha}_{t}}\cdot\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\right)\,+\sigma_{t}\epsilon,$
@@
In a posterior sampling problem, we aim at sampling from the posterior $p(\mathbf{x}|\mathbf{y})$. For this purpose, many recent methods follow the approach described above, using the posterior score $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}|\mathbf{y})=\nabla_{\mathbf{x}}\log p(\mathbf{y}|\mathbf{x}_{t}=\mathbf{x})+\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ instead of $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$. The second term is equivalent to the unconditional score of the pretrained diffusion model, but the challenge lies in the first, i.e., the conditional score term. The conditional probability $p(\mathbf{y}|\mathbf{x}_{t})$ can be written as a conditional expectation $\mathbb{E}_{\mathbf{x}_{0}\sim p(\mathbf{x}_{0}|\mathbf{x}_{t})}[p(\mathbf{y}|\mathbf{x}_{0})]$, but approximating this expectation with Monte Carlo samples is computationally intractable (see Chung et al. [7], for example). For this reason, existing diffusion posterior sampling methods approximate this conditional distribution with a Dirac delta distribution concentrated on $\mathbf{x}_{t}$ [19] or on $\mathbb{E}_{\mathbf{x}_{0}\sim p(\mathbf{x}_{0}|\mathbf{x}_{t})}[\mathbf{x}_{0}]$ [7]. A number of other diffusion posterior sample methods have been proposed [9], each providing a different approximation for the conditional expectation. Most recently, Zhang et al. [43] introduced DAPS (Decoupled Annealing Posterior Sampling), a two-step iterative approach that mitigates the accumulation of errors along the sampling trajectory through a Markov chain Monte Carlo–based equilibration step at $t=0$. DAPS demonstrated state-of-the-art results on both linear and nonlinear inverse problems, but remains limited by the necessity to approximate the conditional probability $p(\mathbf{x}_{0}|\mathbf{x}_{t})$ and by the finiteness of the number of MCMC steps.
@@ 3.2 Dual Ascent Diffusion (DDiff)
At its core, the ADMM method iterates over 3 steps. The $\mathbf{x}$-update corresponds to a data matching step (Eq. 3), the $\mathbf{z}$-update can be seen as a denoising step (Eq. 5) and the dual update stems from the introduction of the dual variable $\mathbf{u}$ for the constraint $\mathbf{x}=\mathbf{z}$ (Eq. 6). While the $\mathbf{x}$ and dual updates are straightforwardly derived from the original ADMM framework, our methodological contribution primarily consists of showing that a pretrained diffusion model can be used more efficiently in the $\mathbf{z}$-update than the naive approach (Eq. 9).
@@
Combined iteratively, these three steps define our method (Algorithm 1). We show that the DDiff updates converge to a fixed point $(\mathbf{x}^{\star},\mathbf{z}^{\star},\mathbf{u}^{\star})$: under bounded denoiser and gradient assumptions, all three iterate sequences are Cauchy and converge to a limit (Theorem 1); notably, this result does not require a convex prior. Full assumptions, proofs, and technical details are in Appendix A.
3.3 DDiff with Latent Diffusion Model
Given a pretrained encoder $\mathcal{E}$ and decoder $\mathcal{D}$, latent diffusion models (LDMs) [29] are trained to model the distribution $p(\mathbf{z}_{0})$ of latent variables $\mathbf{z}_{0}=\mathcal{E}(\mathbf{x}_{0})$, enabling reconstruction via $\mathbf{x}_{0}=\mathcal{D}(\mathbf{z}_{0})$. Leveraging latent diffusion models typically offers reduced memory and compute demands compared to pixel-space diffusion. Extending DDiff to this latent domain yields LatentDDiff, which performs alternating data-fidelity and denoising updates on compact latent representations while preserving the dual-ascent structure of DDiff. The full algorithmic formulation, quantitative results, and discussion of enforcing data consistency either in latent or pixel space are provided in Appendix C.
@@ 3.4 Comparison to Other Diffusion-based Variable Splitting Methods
Prior methods, such as DiffPIR [46], DCDP [24], and PnP-DM [42], are MAP-based optimization methods, specifically built upon the half-quadratic splitting (HQS) method [16], which lacks dual variables in its formulation. Our work naturally extends existing frameworks, transforming them from HQS-style approaches to those using dual variables, including ADMM. By incorporating Lagrange multipliers that accumulate constraint violations across the iterations, DDiff offers improved empirical performance on challenging inverse problems where measurement consistency is crucial. In particular, we note that removing the dual update from DDiff would exactly emulate DiffPIR [46] (with the right choice of $\sigma_{t}$ and $\gamma_{t}$, see Appendix E).
@@
2:for t=T1t=T-1 to 0 do
4:   $\triangleright$ Denoising step (Eq. 11)
5:  $\mathbf{x}\leftarrow\mathbf{z}-\mathbf{u}-\gamma_{t}\nabla_{\mathbf{v}=\mathbf{z}-\mathbf{u}}\|\mathbf{y}-\mathcal{A}(\mathbf{v})\|^{2}$
6:  $\triangleright$ Measurement step (Eq. 10)
7:  $\hat{\epsilon}\leftarrow\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\left(\mathbf{x}_{t}-\sqrt{\bar{\alpha}_{t}}\cdot\mathbf{x}\right)$
8:  $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})\>\textbf{if}\>t>t_{0}\>\textbf{else}\>\epsilon=0$
9:  $\mathbf{x}_{t-1}\leftarrow\sqrt{\bar{\alpha}_{t-1}}\cdot\mathbf{x}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\hat{\epsilon}+\sigma_{t}\epsilon+\sqrt{\bar{\alpha}_{t-1}}\cdot\mathbf{u}$ $\triangleright$ Reverse diffusion (Eq. 12)
10:  $\mathbf{u}\leftarrow\mathbf{u}+\mathbf{x}-\mathbf{z}$ $\triangleright$ Dual update (Eq. 6)
11:end for
12:return $\mathbf{x}_{0}$